# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ Abstract
Recent methodological developments have introduced new black-box approaches to better estimate heterogeneous treatment effects; however, these methods fall short of providing interpretable characterizations of the underlying individuals who may be most at risk or benefit most from receiving the treatment, thereby limiting their practical utility. In this work, we introduce a novel method, causal distillation trees (CDT), to estimate interpretable subgroups. CDT allows researchers to fit any machine learning model of their choice to estimate the individual-level treatment effect, and then leverages a simple, second-stage tree-based model to “distill” the estimated treatment effect into meaningful subgroups. As a result, CDT inherits the improvements in predictive performance from black-box machine learning models while preserving the interpretability of a simple decision tree. We derive theoretical guarantees for the consistency of the estimated subgroups using CDT, and introduce stability-driven diagnostics for researchers to evaluate the quality of the estimated subgroups. We illustrate our proposed method on a randomized controlled trial of antiretroviral treatment for HIV from the AIDS Clinical Trials Group Study 175 and show that CDT out-performs state-of-the-art approaches in constructing stable, clinically relevant subgroups.
11footnotetext: The authors would like to thank participants at the American Causal Inference Conference, Johns Hopkins Biostatistics Causal Inference Seminar, UNC Chapel Hill Biostatistics Seminar for their feedback and comments. 22footnotetext: Denotes equal contribution.33footnotetext: Yale University. Email: melody.huang@yale.edu, URL: www.melodyyhuang.com44footnotetext: University of Notre Dame. Email: ttang4@nd.edu, URL: https://tiffanymtang.github.io/55footnotetext: University of California Irvine. Email: anamaria.kenney@uci.edu
@@ 1 Introduction
While much of the causal inference literature has focused on estimating an average treatment effect for a specific intervention, researchers are often interested in understanding the underlying treatment effect heterogeneity for a given intervention. For example, in medical settings, researchers are often concerned about potential subsets of units who may be harmed by a particular drug or medical intervention. Similarly, in the social sciences, being able to understand who will benefit or could be potentially harmed by a specific intervention is crucial for evaluating the cost-benefit associated with a specific intervention. While different approaches have been proposed to consider how to better estimate treatment effect heterogeneity across individuals, being able to characterize subgroups of individuals in a study in an interpretable way remains challenging.
@@
In its simplest form, many existing subgroup analyses in causal inference focus on subgroups a priori specified by researchers, which relies on substantive knowledge. In the absence of such prior knowledge, many tree-based approaches have been proposed to discover new subgroups with heterogeneous treatment effects (e.g., Su et al., 2009; Lipkovich et al., 2011; Loh et al., 2015; Seibold et al., 2016), of which causal trees (Athey and Imbens, 2016) are arguably the most popular. [^1] Citation count was used as a proxy for popularity. In comparing the citation count of the different papers, as of 2025, the original causal tree paper by Athey and Imbens (2016) has over 2,000 citations, while the alternative methods have anywhere from 100-500 citations. Similar to decision trees (Breiman et al., 1984), causal trees partition the underlying into groups by finding clusters of individuals who have similar treatment effects together. Moreover, unlike linear regression approaches, causal trees output a readily-interpretable partition of subgroups while also accounting for potential high-order interactions.
@@
A drawback to causal trees, however, is that they can be highly unstable and overfit in the presence of even small amounts of noise (e.g., Cattaneo et al., 2022). Recent literature has proposed alternatives to better model heterogeneous treatment effects with ‘metalearners’—e.g., $T$-learners (e.g., Foster et al., 2011), $S$-learners (e.g., Hill, 2011), $X$-learner (Künzel et al., 2019), $R$-learner (Nie and Wager, 2021), $B$-learner (Oprescu et al., 2023), $DR$-learner (Kennedy, 2023), $lp$-$R$-learner (Kennedy et al., 2022). These black-box metalearners provide theoretical guarantees in the form of quasi-oracle properties in recovering the individual-level treatment effect; however, they no longer produce a tree-like output and thus cannot be directly used to identify interpretable subgroups.
@@
In the following paper, we propose causal distillation trees (CDT) to stably estimate interpretable subgroups. CDT allows researchers to leverage the power of black-box meta learners, while preserving the interpretability of a simple tree output. At a high level, CDT is a two-stage learner. The first stage estimates a heterogeneous treatment effect model using a flexible metalearner to generate a predicted individual-level treatment effect for each unit. The second stage then distills the information from the predicted individual-level treatment effect through a decision tree to estimate interpretable subgroups. The first step serves as a de-noising step to help mitigate the instabilities that decision trees traditionally suffer from. The second step of ‘distillation’ draws upon existing ideas in machine learning (Hinton, 2015; Furlanello et al., 2018; Menon et al., 2021), which uses the distillation step to improve prediction accuracy in certain settings. However, unlike typical machine learning settings, where the primary goal is to improve the second stage learner’s predictive ability, we leverage distillation to improve our ability to estimate interpretable subgroups using the second stage decision tree learner.
@@
Our paper provides several key theoretical and methodological contributions. First, while recent papers have similarly proposed the use of a two-stage learner to estimate subgroups (e.g., Foster et al., 2011; Bargagli Stoffi and Gnecco, 2020; Zhang et al., 2021; Rehill, 2024), CDT is agnostic to what first-stage learner researchers use, allowing for a large degree of flexibility in choosing an informative teacher model. Second, unlike previous work, which has largely relied on simulated evidence, we derive theoretical guarantees, proving the consistency of CDT in recovering the optimal partitioning into subgroups. To our knowledge, we are the first paper to prove the theoretical properties in recovering subgroups from using distillation. Our theoretical results highlight that distillation can actually offset the instabilities incurred from using causal trees (or other tree-based approaches) in low signal-to-noise regimes, while preserving the interpretability of these methods.
@@
The paper is structured as follows. Section 2 presents notation and formalizes the notion of a subgroup in causal inference. In Section 3, we introduce causal distillation trees and derive the theoretical properties of CDT. In Section 4, we propose a model selection procedure for researchers to select a teacher model that will maximize stability of subgroup recovery.
We demonstrate the effectiveness of CDT across extensive simulation settings (Section 5), and an AIDS clinical trial case study (Section 6). Section 7 concludes.
@@ 2 Setup and Notation
Throughout the paper, we will use the potential outcomes framework, where $Y_{i}(1)$, $Y_{i}(0)\in\mathbbm{R}$ denote the potential outcome under treatment and control, respectively. Let $Z_{i}\in\{0,1\}$ denote the treatment assignment indicator. We assume consistency of treatment assignment across the study and no interference (i.e., the stable unit treatment value assumption). Let $Y_{i}:=Y_{i}(1)Z_{i}+Y_{i}(0)(1-Z_{i})$ be the observed outcomes. Finally, we assume that researchers have access to a set of pre-treatment covariates $X_{i}\in\mathbbm{R}^{p}$ for each unit in their study, and $\{Y_{i}(1),Y_{i}(0),X_{i},Z_{i}\}$ are drawn independently and identically distributed from an arbitrary joint distribution for $i=1,...,n$.
A common estimand of interest is the average treatment effect for a given study sample:
@@
In practice, a more policy-relevant quantity is the conditional average treatment effect, which considers the average treatment effect across subsets of individuals in the study population. Existing literature has recommended researchers leverage their substantive expertise to posit what characteristics could potentially moderate the treatment effect.
Informally, we consider a subgroup as a subset of individuals, defined by a discrete partition in the space of $X$, where the partitions in the space correspond to substantively different treatment effects. In particular, we are interested in characterizing different groups of individuals in a study that have different effects from receiving the same treatment.
We formally define a subgroup $\mathcal{G}_{g}:\mathcal{X}\to\{0,1\}$ as
$\mathcal{G}_{g}(X)=\prod_{j=1}^{p}\mathbbm{1}\left\{X^{(j)}\in R^{(j)}_{g}\right\},$ (1)
where $X^{(j)}$ denotes the $j^{th}$ covariate, and $R^{(j)}_{g}\subseteq\text{supp}(X^{(j)})$. Here, $\mathcal{G}_{g}(X)$ represents a collection of binary decision rules that characterize a subset of individuals in the study (Lipkovich et al., 2017). For example, if $\mathcal{G}_{g}(X)=\mathbbm{1}\{X^{\text{gender}}=\text{Female}\}\cdot\mathbbm{1}\{X^{\text{age}}<35\}$, then $\mathcal{G}_{g}(X)$ corresponds to the subset of female-identifying individuals who are under the age of 35. A collection of subgroups $\{\mathcal{G}_{1}(X_{i}),\ldots,\mathcal{G}_{G}(X_{i})\}$ is said to be a discrete partition of the covariate space of $X$ if $\sum_{g=1}^{G}\mathcal{G}_{g}(X_{i})=1$ for all units $i=1,\ldots,n$. (i.e., each unit belongs to exactly one subgroup and cannot belong to multiple subgroups).
Finally, given a subgroup $\mathcal{G}_{g}(X)$, we define $\tau^{(g)}$ as the subgroup average treatment effect:
$\tau^{(g)}:=\frac{1}{n_{g}}\sum_{i=1}^{n}\mathcal{G}_{g}(X_{i})\cdot\left\{Y_{i}(1)-Y_{i}(0)\right\},$ (2)
where $n_{g}$ is the number of samples in subgroup $g$.
The primary focus of the paper will be on estimating an optimal partition, which maximizes treatment effect heterogeneity across subgroups using the most parsimonious set of rules. Specifically, we aim to construct the most parsimonious partition that minimizes the squared loss between the true individual-level treatment effects and the associated subgroup average treatment effects:
@@
where $\Omega$ denotes the space of all possible partitions of the covariate space of $X$. Under squared loss, the optimal partition corresponds to a subset in the covariate space of $X$ that maximizes the difference in treatment effect heterogeneity across subgroups. In other words, within a partitioned group, the treatment effect heterogeneity is minimized, thereby maximizing the variance in treatment effects across groups. In practice, researchers can choose alternative loss functions that correspond to what are substantively meaningful partitions in the covariate space. However, the theoretical guarantees we derive utilize the squared loss, and generalizing the results for other loss functions is an open avenue of future research.
Our work is distinct from existing literature in considering subgroup average treatment effects. For example, alternative quantities considered in the literature include the sorted group average treatment effect (GATES), which considers the average treatment effect, across subsets formed based on the magnitude of the estimated effect (i.e., the ATE, across individuals who would most benefit) (e.g., Imai and Li, 2024; Chernozhukov et al., 2018; Dwivedi et al., 2020). GATES relies on using a machine learning model to generate predictions of the CATE. Then, grouping the predicted CATEs by highest to lowest, it estimates the ATE across each group. While quantities like GATE provide a way for researchers to evaluate the range of possible treatment effects within a study, it falls short of providing interpretability for the characteristics of the units within each group.
@@ 3.1 Overview
We propose a method, causal distillation trees (CDT), which stably estimates interpretable subgroups. CDT is a two-stage learner (Figure 1). The first stage learner, referred to as the teacher model, learns an informative model for the individual-level treatment effect using the observed data $\{Y,Z,X\}$. Using this teacher model, we generate a prediction for the individual-level treatment effect for all units (i.e., $\hat{\tau}(X_{i})$). In the second stage, we distill the predicted $\hat{\tau}(X_{i})$’s into interpretable subgroups by fitting a decision tree, called the student model, to predict the $\hat{\tau}(X_{i})$’s from the covariates $X$. The estimated decision rules from the decision tree define a discrete partition, which naturally maps to the collection of subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$ (see Figure 2 for illustration). Finally, using these estimated partitions, we estimate the subgroup average treatment effect for each subgroup.
@@
Figure 1: Overview of Causal Distillation Trees (CDT). CDT leverages a two-stage procedure, which first fits a teacher model (e.g., a black-box metalearner) to estimate individual-level treatment effects and secondly fits a student model (e.g., a decision tree) to predict the estimated individual-level treatment effects, in effect distilling the estimated individual-level treatment effects and producing interpretable subgroups. This two-stage learner is learned using the training data (blue-gray boxes). Finally, using the estimated subgroups, the subgroup average treatment effects are honestly estimated with a held-out estimation set (yellow box).
@@
Informally, we can think of the first stage learner as smoothing the individual-level treatment effects by projecting $\tau_{i}$ into the basis space of the covariates $X$. By using the projected version of the individual-level treatment effect instead of $\tau_{i}$, the second stage learner will be able to more stably estimate the subgroups, as the teacher model will have de-noised the outcomes.
@@
Throughout the paper, we use a decision tree—specifically, a classification and regression tree (CART) (Breiman et al., 1984)—as the second-stage learner. While researchers can in theory swap out the second-stage learner for alternative rule-based approaches (e.g., Bargagli-Stoffi et al., 2020; Wang and Rudin, 2022; Wan et al., 2023), we focus on decision trees because the output of the decision tree maps intuitively to a standard interpretation of a subgroup as a discrete partition of the covariate space.
@@
To avoid post-selection bias, we recommend researchers use sample splitting. Specifically, researchers should randomly split the data into a training and a hold-out estimation set. The training set will be used to estimate the subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$ while the hold-out estimation set is reserved to estimate the subgroup average treatment effects $\hat{\tau}^{(1)},\ldots,\hat{\tau}^{(G)}$. This provides an honest estimate of the subgroup average treatment effect (Athey and Imbens, 2016).
In addition, to avoid overfitting to the teacher model, researchers should estimate the $\hat{\tau}(X_{i})$’s using out-of-sample procedures in the first stage. In particular, if the teacher model comes with a built-in out-of-sample estimation procedure (e.g., causal forests can estimate $\hat{\tau}(X_{i})$ via out-of-bag samples), we estimate $\hat{\tau}(X_{i})$ using these model-specific out-of-sample estimates. Otherwise, we perform repeated cross-fitting (Chernozhukov et al., 2018) to learn the $\hat{\tau}(X_{i})$’s. We detail this sample splitting procedure and CDT in Algorithm 2.
Figure 2: Given a decision tree (A), its decision splits correspond to binary cuts that partition the covariate space $\mathcal{X}$, shown in (B). Furthermore, the decision splits leading to each terminal node map to the collection of rules that define a subgroup, shown in (C).

### v2

@@ Abstract
Recent methodological developments have introduced new black-box approaches to better estimate heterogeneous treatment effects; however, these methods fall short of providing interpretable characterizations of the underlying individuals who may be most at risk or benefit most from receiving the treatment, thereby limiting their practical utility. In this work, we introduce causal distillation trees (CDT) to estimate interpretable subgroups. CDT allows researchers to fit any machine learning model to estimate the individual-level treatment effect, and then leverages a simple, second-stage tree-based model to “distill” the estimated treatment effect into meaningful subgroups. As a result, CDT inherits the improvements in predictive performance from black-box machine learning models while preserving the interpretability of a simple decision tree. We derive theoretical guarantees for the consistency of the estimated subgroups using CDT, and introduce stability-driven diagnostics for researchers to evaluate the quality of the estimated subgroups. We illustrate our proposed method on a randomized controlled trial of antiretroviral treatment for HIV from the AIDS Clinical Trials Group Study 175 and show that CDT out-performs state-of-the-art approaches in constructing stable, clinically relevant subgroups.
11footnotetext: The authors would like to thank Sam Pimentel, Yan Shuo Tan, and participants at the American Causal Inference Conference, Johns Hopkins Biostatistics Causal Inference Seminar, UNC Chapel Hill Biostatistics Seminar for their feedback and comments.22footnotetext: Denotes equal contribution.33footnotetext: Yale University. Email: melody.huang@yale.edu, URL: www.melodyyhuang.com44footnotetext: University of Notre Dame. Email: ttang4@nd.edu, URL: https://tiffanymtang.github.io/55footnotetext: University of California Irvine. Email: anamaria.kenney@uci.edu
Keywords: causal inference, treatment effect heterogeneity, subgroup estimation
@@ 1 Introduction
While much of the causal inference literature has focused on estimating an average treatment effect for a specific intervention, researchers are often interested in understanding the underlying treatment effect heterogeneity for a given intervention. For example, in medical settings, researchers are often concerned about potential subsets of units who may be harmed by a particular drug or medical intervention. Similarly, in the social sciences, being able to understand who will benefit or could be potentially harmed by a specific intervention is crucial for evaluating the cost-benefit associated with a specific intervention. Though different approaches have been proposed to consider how to better estimate treatment effect heterogeneity across individuals, being able to characterize subgroups of individuals in a study in an interpretable way remains challenging.
@@
In its simplest form, many subgroup analyses in causal inference focus on pre-specified subgroups based upon prior substantive knowledge. However, it can still be challenging to reason about higher-order interactions between covariates. In the absence of prior knowledge, different tree-based approaches have been proposed to discover new subgroups with heterogeneous treatment effects (e.g., Su et al., 2009; Lipkovich et al., 2011; Loh et al., 2015), of which causal trees (Athey & Imbens, 2016) are arguably the most popular. [^1] Citation count was used as a proxy for popularity. In comparing the citation count of the different papers, as of 2025, the original causal tree paper by Athey & Imbens (2016) has over 2,000 citations, while the alternative methods have anywhere from 100-500 citations. Similar to decision trees (Breiman et al., 1984), causal trees partition the underlying covariate space into groups by finding clusters of individuals who have similar treatment effects. Moreover, unlike linear regression approaches, causal trees output a readily-interpretable partition of subgroups while also accounting for potential high-order interactions.
@@
Unfortunately, a drawback to causal trees is that they can be highly unstable and overfit in the presence of even small amounts of noise (e.g., Cattaneo et al., 2022). Recent literature has proposed alternatives to better model heterogeneous treatment effects with ‘metalearners’—e.g., $T$-learners (e.g., Foster et al., 2011), $S$-learners (e.g., Hill, 2011), $X$-learner (Künzel et al., 2019), $R$-learner (Nie & Wager, 2021), to name a few. These black-box metalearners provide theoretical guarantees in the form of quasi-oracle properties in recovering the individual-level treatment effect; however, they no longer produce a tree-like output and thus cannot be directly used to identify interpretable subgroups.
@@
In the following paper, we propose causal distillation trees (CDT) to stably estimate interpretable subgroups. CDT allows researchers to leverage the power of black-box meta learners, while preserving the interpretability of a simple tree output. CDT is a two-stage learner. The first stage estimates a heterogeneous treatment effect model using a flexible metalearner to generate a predicted individual-level treatment effect for each unit. The second stage then distills the information from the predicted individual-level treatment effect through a decision tree to construct interpretable subgroups. The first step serves as a de-noising step to help mitigate the instabilities that decision trees traditionally suffer from. The second step of ‘distillation’ draws upon existing ideas in machine learning (e.g., Hinton, 2015; Menon et al., 2021; Dao et al., 2021). However, unlike typical machine learning settings, where the primary goal is to improve the second stage learner’s predictive ability, we leverage distillation to improve our ability to estimate interpretable subgroups using the second stage decision tree learner.
@@
Our paper provides several key theoretical and methodological contributions. First, while recent papers have similarly proposed the use of a two-stage learner to estimate subgroups (e.g., Foster et al., 2011; Zhang et al., 2021; Rehill, 2024), CDT is agnostic to what first-stage learner researchers use, allowing for a large degree of flexibility in choosing an informative teacher model. Second, unlike previous work, which has largely relied on simulated evidence, we derive theoretical guarantees, proving that CDT consistently recovers the optimal subgroups. To our knowledge, we are the first paper to prove the theoretical properties in recovering subgroups from distillation. Our theoretical results highlight that distillation can offset the instabilities incurred from using causal trees (or other tree-based approaches) in low signal-to-noise regimes, while preserving the interpretability of these methods.
@@
The paper is structured as follows. Section 2 presents notation and formalizes the notion of a subgroup in causal inference. In Section 3, we introduce causal distillation trees and derive its theoretical properties. In Section 4, we propose a stability-driven model selection procedure to select a teacher model in CDT. We demonstrate the effectiveness of CDT across extensive simulation settings (Section 5), and an AIDS clinical trial case study (Section 6). Section 7 concludes.
@@ 2 Setup and Notation
[…]
A common estimand of interest is the average treatment effect (ATE) for a given study:
@@
In practice, a more policy-relevant quantity is the conditional average treatment effect (CATE), which considers the ATE across subsets of individuals in the study population. Existing literature has recommended researchers leverage their substantive expertise to posit what characteristics could potentially moderate the treatment effect.
The focus of this paper is to consider how to construct interesting subgroups for subsets of individuals. We consider a subgroup as a subset of individuals, defined by a discrete partition in the space of $X$, where the partitions in the space correspond to substantively different treatment effects. In particular, we are interested in characterizing different groups of individuals in a study that have different effects from receiving the same treatment.
We formally define a subgroup $\mathcal{G}^{\prime}_{g}:\mathcal{X}\to\{0,1\}$ as
$\mathcal{G}^{\prime}_{g}(X)=\prod_{j=1}^{p}\mathbbm{1}\left\{X^{(j)}\in R^{(j)}_{g}\right\},$ (1)
where $X^{(j)}$ denotes the $j^{th}$ covariate, and $R^{(j)}_{g}\subseteq\text{supp}(X^{(j)})$. Here, $\mathcal{G}^{\prime}_{g}(X)$ represents a collection of binary decision rules that characterize a subset of individuals in the study (Lipkovich et al., 2017). For example, if $\mathcal{G}^{\prime}_{g}(X)=\mathbbm{1}\{X^{\text{gender}}=\text{Female}\}\cdot\mathbbm{1}\{X^{\text{age}}<35\}$, then $\mathcal{G}^{\prime}_{g}(X)$ corresponds to the subset of female-identifying individuals who are under the age of 35. Throughout the paper, we focus on subgroups that are a discrete partition of the covariate space–i.e., $\sum_{g=1}^{G}\mathcal{G}^{\prime}_{g}(X_{i})=1$ for all units $i=1,\ldots,n$ (i.e., each unit belongs to exactly one subgroup).
Given a subgroup $\mathcal{G}^{\prime}_{g}(X)$, we define $\tau^{(g)}$ as the subgroup average treatment effect:
$\tau^{(g)}:=\mathbb{E}\left[Y(1)-Y(0)\mid\mathcal{G}^{\prime}_{g}(X)=1\right].$ (2)
The primary aim in this paper is to estimate an optimal partition, which maximizes treatment effect heterogeneity across subgroups using the most parsimonious set of rules. Specifically, we aim to construct the most parsimonious partition that minimizes the squared loss between the true individual-level treatment effects and the associated subgroup ATEs:
@@
where $\Omega$ represents a class of partitions in the covariate space of $X$. Under squared loss, the optimal partition corresponds to a subset in the covariate space of $X$ that maximizes the difference in treatment effect heterogeneity across subgroups. In other words, within a partitioned group, the treatment effect heterogeneity is minimized, thereby maximizing the variance in treatment effects across groups. In practice, researchers can choose alternative loss functions that correspond to what are substantively meaningful partitions in the covariate space. However, the theoretical guarantees we derive utilize the squared loss, and generalizing the results for other loss functions is an open avenue of future research.
Our work is distinct from existing literature in considering subgroup ATEs. For example, alternative quantities in the literature include the sorted group average treatment effect (GATES), which considers the ATE across subsets formed based on the magnitude of the estimated effect (i.e., the ATE across individuals who would most benefit) (e.g., Imai & Li, 2024; Chernozhukov, Demirer, Duflo & Fernandez-Val, 2018; Dwivedi et al., 2020). GATES relies on using a machine learning model to generate predictions of the CATE, and then groups the predicted CATEs by highest to lowest to estimates the ATE across each group. While quantities like GATE provide a way for researchers to evaluate the range of possible treatment effects within a study, it falls short of providing interpretability for the characteristics of the units within each group.
@@ 3.1 Overview
We propose a method, causal distillation trees (CDT), which stably estimates interpretable subgroups. CDT is a two-stage learner (Figure 1 and Algorithm 1). The first stage learner, referred to as the teacher model, learns an informative model for the individual-level treatment effect using the observed data $\{Y,\;Z,\;X\}$. Using this teacher model, we generate a prediction $\hat{\tau}_{i}^{d}$ for the individual-level treatment effect for all units. In the second stage, we distill the predicted $\hat{\tau}_{i}^{d}$’s into interpretable subgroups by fitting a decision tree, called the student model, to predict the $\hat{\tau}_{i}^{d}$’s from the covariates $X$. The estimated decision rules from the decision tree define a discrete partition, which naturally maps to the collection of subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\;\ldots,\;\hat{\mathcal{G}}_{G}(X)\right\}$ (see Figure A1). Finally, using the estimated partitions, we estimate the subgroup ATE for each subgroup.
@@
Figure 1: Overview of Causal Distillation Trees (CDT). CDT leverages a two-stage procedure, which first fits a teacher model to estimate individual-level treatment effects and then fits a student model (e.g., a decision tree) to distill the estimated individual-level treatment effects and produce interpretable subgroups. The two-stage learner is learned using the training data (blue-gray boxes). Using the estimated subgroups, the subgroup ATEs are honestly estimated with a held-out estimation set (yellow box).
@@
Informally, the first stage learner smooths the individual-level treatment effects by projecting $\tau_{i}$ into the basis space of the covariates $X$. By using the projected version of the individual-level treatment effect instead of $\tau_{i}$, the second stage learner will be able to more stably estimate the subgroups, as the teacher model will have de-noised the outcomes.
@@
Throughout the paper, we use a decision tree—specifically, a classification and regression tree (CART) (Breiman et al., 1984)—as the second-stage learner. While researchers can use alternative models as the second-stage learner (e.g., Bargagli-Stoffi et al., 2020; Wang & Rudin, 2022; Wan et al., 2023), we focus on decision trees because the output maps intuitively to a standard interpretation of a subgroup as a discrete partition of the covariate space. In settings when researchers are interested in a linear CATE, it would be useful to consider alternative second-stage learners.
@@
We also incorporate careful sample splitting in CDT to avoid post-selection bias. Specifically, we randomly split the data into a training set to estimate the subgroups $\left\{\hat{\mathcal{G}}_{1}(X),\ldots,\hat{\mathcal{G}}_{G}(X)\right\}$ and a hold-out estimation set to estimate the subgroup ATEs $\hat{\tau}^{(1)},\ldots,\hat{\tau}^{(G)}$ (see Figure 1). This provides an honest estimate of the subgroup ATE (Athey & Imbens, 2016). In addition, to avoid overfitting to the teacher model, we estimate the $\hat{\tau}_{i}^{d}$’s using out-of-sample procedures (e.g., using out-of-bag samples (Wager & Athey, 2018) or cross-fitting (Chernozhukov, Chetverikov, Demirer, Duflo, Hansen, Newey & Robins, 2018)) in the first stage (details in Algorithm 1).